OpenStudy (anonymous):

A laboratory has 20 grams of a radioactive substance that decays in such a way that t days later the number of grams left will be y = 20 (1/10)^(t/15). How many days does it take for half of the substance to decay?

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

half of 20 is 10 .... set it equal to 10 and solve for "t"

OpenStudy (dumbcow):

remember \[\ln (a^{t}) = t*\ln a\]

OpenStudy (anonymous):

ok can you write out the equation for me i still dont get it much

OpenStudy (dumbcow):

\[20*(1/10)^{t/15} = 10\] \[(1/10)^{t/15} = 1/2\] take log of both sides

OpenStudy (anonymous):

log(1/2)/ log 1/10 i got .30.....

OpenStudy (dumbcow):

yes .301

OpenStudy (anonymous):

thn multiply 15 right

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

4.5 Thanks :D

OpenStudy (dumbcow):

:)